OpenStudy (anonymous):

cos^2-sin(2theta)= sin^2 theta

OpenStudy (anonymous):

?

OpenStudy (klimenkov):

\(\cos^2\theta-\sin2\theta=\sin^2\theta\)

OpenStudy (anonymous):

Try:\[\cos^{2}(\theta)=1-\sin ^{2}(\theta)\] on \[\cos^{2}(\theta)\] and simplify.

OpenStudy (anonymous):

Gudnyt.

OpenStudy (anonymous):

hey but \[\sin ^{2}+\cos ^{2}=1 \] so like i get stuck where 1=2sinthetacostheta